# Differentially Private Next-Token Prediction of Large Language Models

### v1

@@ Abstract
Ensuring the privacy of Large Language Models (LLMs) is becoming increasingly important. The most widely adopted technique to accomplish this is DP-SGD, which trains a model in such a way that guarantees Differential Privacy (DP). However, DP-SGD requires longer training times and larger memory requirements than SGD, while overestimating an adversary’s capabilities in having white box access to the model. A more realistic scenario assumes only black-box access to a privacy-sensitive LLM. Motivated by these observations, we present Private Mixing of Ensemble Distributions (PMixED): a private prediction protocol that achieves practical next-token prediction by projecting each of the model’s output distribution from an ensemble of fine-tuned LLMs onto a set around a public LLM’s output distribution, then averaging the projected distributions and sampling from it. Our approach is more lightweight than DP-SGD in that it is model agnostic, instead providing differential privacy at prediction rather than during training. Our results show that PMixED achieves a stronger privacy guarantee than sample-level privacy and outperforms DP-SGD for privacy $\epsilon=8$ on large-scale datasets.
@@ 1 Introduction
Large language models (LLMs) are being deployed in many facets of our lives, from email autocomplete tools to interactive chatbots. Their commercial success is largely attributed to their ability to generate human-like text. However, when given query access to an LLM, it has been shown that LLMs are susceptible to training data extraction attacks Carlini et al. (2021) due to memorization of training samples Carlini et al. (2019). These security vulnerabilities have recently catalyzed government intervention, most notably the EU’s AI Act EU (2024) and the US executive order on Safe, Secure, and Trustworthy AI US (2023). Thus, it is becoming a requirement that entities deploying LLMs must do so in a privacy-preserving way.
The gold standard for achieving strong privacy is Differential Privacy (DP), which is a mathematical framework that reduces the influence each data sample has on the overall output of an algorithm Dwork (2006). The most widely known technique injects DP noise into the training algorithm, called DP-SGD Abadi et al. (2016).
Recent works have shown that applying DP-SGD during fine-tuning on pre-trained LLMs provides acceptable results Li et al. (2021); Yu et al. (2021). Unfortunately, scaling these results to larger datasets and models becomes challenging since (1) The magnitude of the noise added by DP-SGD scales with the total number of parameters of the model, i.e., $\sqrt{d}$ Kamath (2020), in the worst case; and (2) ML accelerated hardware is designed to exploit batch operations but is underutilized by DP-SGD since it requires per-sample gradient calculations, which cause large runtime and memory consumption Yousefpour et al. (2021). Most importantly, many commercial deployments of LLMs are only accessible via an API, e.g. Chat-GPT, yet DP-SGD assumes an adversary has access to the model parameters, which can lead to overestimating the privacy loss bounds Nasr et al. (2021).
@@
To address the aforementioned limitations, we propose PMixED, depicted in Figure 1. Rather than providing DP during training, PMixED instead provides DP for a private corpus during inference by utilizing a public model [^1] It is assumed that a public LLM does not leak the privacy of a private corpus. However, in practice, it can, although works in differentially private fine-tuning of a public model implicitly use this assumption. to mix the predictions of non-private LLMs. This approach can be broken down into two phases: training and private prediction. During Phase-I: training, PMixED follows the sample-and-aggregate paradigm Nissim et al. (2007) by partitioning a private corpus into $N$ pairwise disjoint subsets, then fine-tuning a pre-trained LLM on each subset to produce an ensemble. Using an ensemble is crucial for guaranteeing privacy, but also has been shown to boost perplexity Jozefowicz et al. (2016). During Phase-II: private prediction, PMixED selects a subset of the ensemble using Poisson sampling. Each selected model produces its output probability distribution and projects it onto a set around a public model’s output distribution. Lastly, PMixED averages these optimal projections and then samples from it.
Because PMixED does not employ differentially private training algorithms, it does not incur substantial training overhead like DP-SGD. Furthermore, we reduce computational and storage costs by employing parameter-efficient fine-tuning methods. In summary, our key contributions are the following: (1) We introduce $(\alpha,\beta)$-RD Mollifiers which generalize $\epsilon$-Mollifers introduced in Husain et al. (2020) and avoid additive DP noise. (2) We propose a private prediction protocol utilizing RD Mollifiers and prove that it satisfies the DP prediction definition with group-level privacy. (3) We experimentally demonstrate that PMixED outperforms DP-SGD on two large-scale datasets. (4) We open-source our software implementation of PMixED to further spark research in this area [^2] https://github.com/james-flemings/pmixed.
@@ 2 Related Works
Most of the differential privacy work in LLMs has focused on private training. McMahan et al. (2017) first explored private training of language models using a small recurrent neural network to achieve user-level differential privacy in the federated learning setting. Recent breakthroughs in differentially private LLMs involve self-supervised pre-training on public data, followed by privately fine-tuning on a private corpus Li et al. (2021); Yu et al. (2021).
An orthogonal approach, but conceptually similar to ours, is PATE Papernot et al. (2016, 2018). PATE also uses an ensemble of models trained on pairwise disjoint subsets of a private dataset to make DP predictions. However, PATE and PMixED rely on substantially different private aggregation schemes. PATE uses the Gaussian mechanism to perturb its output vote count histogram while our method utilizes the inherent stochastic nature of sampling from a probability distribution and a public model. Furthermore, PATE can not be used for private prediction since its data-dependent privacy accounting causes its privacy loss to be a function of the private data, which leaks additional privacy when testing on it Redberg et al. (2023).
@@ 3.2 Differential Privacy
3.3 Private Training & Prediction
We say that $A$ is a private training algorithm if it produces weights that are differentially private with respect to a private corpus $D$. For DP-SGD, $A(D)$ returns the per-sample gradients of the parameters perturbed with Gaussian noise for each iteration of training. However, in commercial deployments of LLMs, an adversary does not have access to model parameters. An alternative approach, private prediction Dwork and Feldman (2018), exploits this important relaxation by providing DP at prediction. We formalize this below:
@@ 5.1 DP Guarantees for PMixED
At a high level, we will prove that $\mathcal{P}$ is $(\alpha,\epsilon_{G}/T)$-RDP for query $\mathbf{x}_{t}$. Then, using the Composition Theorem C.2, we can say that $\mathcal{P}$ is $(\alpha,\epsilon_{G})$-RDP for query-responses $\{(\mathbf{x}_{t},y_{t})\}_{t=1}^{T}$, thus satisfying the Private Prediction Definition 3.4. Appendix B analyzes the case when $\alpha=\infty$, which is pure DP. Essentially, $D_{\alpha}(p(\mathbf{x}_{t})||p_{-i}(\mathbf{x}_{t}))$ gives an upper bound while $D_{\alpha}(p_{-i}(\mathbf{x}_{t})||p(\mathbf{x}_{t}))$ gives a lower bound for $\beta$ Thus, it suffices to show $D_{\alpha}(p(\mathbf{x}_{t})||p_{-i}(\mathbf{x}_{t}))\leq\epsilon_{G}/T$.
@@
where $\lambda_{i}$ is selected from Equation 5. Now, observe that each $\lambda_{i}$ is dependent only on $D_{i}$, so $p_{-i}(\mathbf{x}_{t})$ does not contain $D_{i}$. Using the fact that $D_{\alpha}^{\leftrightarrow}(p_{j}(\mathbf{x}_{t})||p_{0}(\mathbf{x}_{t}))\leq\beta\alpha$ for all $j\in[N]$, then for any two neighboring ensembles $\{p_{i}\}_{i=1}^{N}$, $\{p_{j}\}_{j\neq i}$ with $N>1$
@@
by the Quasi Convexity property of Renyi Divergence C.6. Then using the Triangle-like Inquality C.5 gives us $D_{\alpha}(\overline{p}_{i}(\mathbf{x}_{t})||p_{-i}(\mathbf{x}_{t}))\leq 4\beta\alpha$. Hence plugging in Equation 6 for $\beta$ into Equation 8 implies $D_{\alpha}(p(\mathbf{x}_{t})||p_{-i}(\mathbf{x}_{t}))\leq\epsilon_{G}/T$. When $N=1$, then $p_{-i}(\mathbf{x}_{t})=p_{0}(\mathbf{x}_{t})$ since our protocol will resort to using $p_{0}$. Hence $D_{\alpha}(p(\mathbf{x}_{t})||p_{-i}(\mathbf{x}_{t}))\leq\beta\alpha\leq\epsilon_{G}/T$. Thus proves the claim that the output of $\mathcal{P}$ on query $\mathbf{x}_{t}$ is $(\alpha,\epsilon_{G}/T)$-RDP. ∎
@@ 6.1 Experimental Setup
The public model in our experiments is a pre-trained GPT-2 small model. We compare PMixED to three baselines: the public model, a non-private fine-tuned model, and a private fine-tuned model produced by DP-SGD. Although DP-SGD has a weaker privacy level, we compared PMixED to per-sample DP-SGD as our baseline because it illustrates how well PMixED performs against the most widely-used DP solution. The non-private and private fine-tuned baselines also used LoRA. Since LoRA uses a much smaller set of parameters that are updated during fine-tuning, it can outperform full private fine-tuning Yu et al. (2021). The LoRA parameters used for each model $p_{i}$ contain $0.11\%$ of the total number of parameters of the pre-trained model, allowing us to fit the entire ensemble into one GPU during prediction. Appendix A contains more details about fine-tuning.
@@ 6.3 Ablation Study
7 Conclusion and Future Directions
In this work, we introduced a novel private prediction protocol that can be used when we only have API access to an LLM. By utilizing the inherit stochasticity of next-token prediction and RD mollifiers to reduce memorization of a private corpus, we can avoid adding DP noise thereby mitigating utility loss. Furthermore, PMixED is compatible with various privacy granularities, being at least stronger than sample-level DP. We experimentally showed that PMixED outperforms DP-SGD in terms of utility on large-scale datasets using LLM. Thus, PMixED substantially progresses private prediction methods in LLMs and offers a practical alternative to DP training methods.
Although our experiments only used a pre-trained GPT-2 small as the public model, in general, we can plug and play any public model for inference, highlighting the versatility of our approach. For example, we could use a GPT-2 model that is already fine-tuned on a public corpus from a different domain, or we could use a larger pre-trained model like GPT-2 XL. As a result, we can further boost the performance of our approach without additional training costs. We leave this exploration as a future work. Lastly, PMixED naturally applies to various privacy levels in the user-level setting, which is another possible direction to explore.
@@ 8 Limitations
PMixED does not require per-sample gradients, thus significantly reducing the training overhead compared to DP-SGD. However, as shown in the ablation studies, PMixED performs better with larger ensemble sizes combined with longer training epochs. Thus, it needs large amounts of data so that each model in the ensemble has enough data to learn, while needing more training epochs than regular SGD. However, both limitations are not unique to PMixED as DP-SGD suffers from this as well.
In this work, we employed LoRA to significantly reduce storage. However, we do acknowledge that storing an ensemble in memory can be cumbersome to add to the training and inference procedure. Moreover, the additional inference latency involved with sequential inference of the ensemble and calculating the lambdas, due to the bisection method, results in some additional overhead. Potential systems optimizations can be to parallelize the inference and lambda calculations. We leave it as a future work to reduce the inference latency.
Perhaps the biggest limitation of PMixED is the query budget, which limits the number of predictions that can be made. However, this limitation is a consequence of differentially private prediction from a nonprivately trained model, as it also appears in other works. As a future work, we are exploring a more relaxed problem setup, a PATE-like setting Papernot et al. (2016, 2018), where we produce DP predictions while minimizing the privacy loss.
Finally, we note how work in DP-SGD, as well as this work, uses a pre-trained model to boost performance. Unfortunately, using publicly available data is not necessarily risk-free in terms of privacy, as prior works were able to extract personally identifiable information from a GPT-2 model pre-trained on data scraped from the public internet Carlini et al. (2021).

### v3

@@ Abstract
Ensuring the privacy of Large Language Models (LLMs) is becoming increasingly important. The most widely adopted technique to accomplish this is DP-SGD, which trains a model to guarantee Differential Privacy (DP). However, DP-SGD overestimates an adversary’s capabilities in having white box access to the model and, as a result, causes longer training times and larger memory usage than SGD. On the other hand, commercial LLM deployments are predominantly cloud-based; hence, adversarial access to LLMs is black-box. Motivated by these observations, we present Private Mixing of Ensemble Distributions (PMixED): a private prediction protocol for next-token prediction that utilizes the inherent stochasticity of next-token sampling and a public model to achieve Differential Privacy. We formalize this by introducing RD-mollifers which project each of the model’s output distribution from an ensemble of fine-tuned LLMs onto a set around a public LLM’s output distribution, then average the projected distributions and sample from it. Unlike DP-SGD which needs to consider the model architecture during training, PMixED is model agnostic, which makes PMixED a very appealing solution for current deployments. Our results show that PMixED achieves a stronger privacy guarantee than sample-level privacy and outperforms DP-SGD for privacy $\epsilon=8$ on large-scale datasets. Thus, PMixED offers a practical alternative to DP training methods for achieving strong generative utility without compromising privacy.
@@ 1 Introduction
Large language models (LLMs) are being deployed to improve societal productivity, from troubleshooting complex systems to autocompletion tools and interactive chatbots. Their commercial success is largely attributed to their ability to generate human-like text. However, when given query access to an LLM, it has been shown that LLMs are susceptible to training data extraction attacks Carlini et al. (2021) due to memorization of training samples Carlini et al. (2019). These security vulnerabilities have recently catalyzed government intervention, most notably the EU’s AI Act EU (2024) and the US executive order on Safe, Secure, and Trustworthy AI US (2023). Thus, it is becoming a requirement that entities deploying LLMs must do so in a privacy-preserving way.
The gold standard for achieving strong privacy is Differential Privacy (DP), a mathematical framework that reduces how much an LLM memorizes individual data samples Dwork (2006). The most widely known technique injects strategic noise into the training algorithm, called DP-SGD Abadi et al. (2016). It has recently been shown that applying DP-SGD during fine-tuning on pre-trained LLMs provides acceptable results Li et al. (2021); Yu et al. (2021). Unfortunately, scaling these results to larger datasets and models becomes challenging since (1) The magnitude of the noise added by DP-SGD scales with the total number of parameters of the model, i.e., $\sqrt{d}$ Kamath (2020), in the worst case; and (2) ML accelerated hardware is designed to exploit batch operations but is underutilized by DP-SGD since it requires per-sample gradient calculations, which cause large runtime and memory consumption Yousefpour et al. (2021).
@@
A key observation that motivates this work is that many commercial deployments of LLMs are only accessible via an API, e.g. Chat-GPT. Hence, an adversary has only black-box access to the model, yet DP-SGD assumes the adversary has white-box access to the model since it applies differential privacy to the model parameters. Such a pessimistic assumption on adversarial capabilities can lead to overestimating the privacy loss bounds Nasr et al. (2021). We believe that improving the privacy of LLMs under the black box assumption is the key to enabling wider adoption of privacy-preserving LLMs. However, prior work has demonstrated that private prediction algorithms generally perform worse than private training algorithms van der Maaten and Hannun (2020).
To address the aforementioned limitations, we propose PMixED, depicted in Figure 1. Rather than providing DP during training, PMixED instead provides DP for a private corpus during inference by utilizing two crucial features: (1) Randomness comes for free when predicting the next-token by sampling from the output probability distribution of a language model; (2) We can bound the privacy leakage of a prediction using a public model [^1] While it is assumed that a public LLM does not compromise the privacy of a private corpus, in practice, such models can inadvertently leak information. This assumption is implicitly relied upon in works that employ differentially private fine-tuning of public models. to mix the predictions of privacy-sensitive LLMs. We formalize these two observations by introducing $(\alpha,\beta)$-RD Mollifiers which generalize $\epsilon$-Mollifers introduced in Husain et al. (2020) and could be of interest independent of this work.
PMixED can be broken down into two phases: training and private prediction. During Phase-I: training, PMixED follows the sample-and-aggregate paradigm Nissim et al. (2007) by partitioning a private corpus into $N$ pairwise disjoint subsets, then fine-tuning a pre-trained LLM on each subset to produce an ensemble. Using an ensemble is crucial for guaranteeing privacy, but also has been shown to boost perplexity Jozefowicz et al. (2016). During Phase-II: private prediction, PMixED selects a subset of the ensemble using Poisson subsampling. Each selected model produces its output probability distribution, then follows the RD-mollification process by projecting its output distribution onto a Renyi Divergence ball centered at a public model’s output distribution. Lastly, PMixED averages these optimal projections and then samples from it.
Because PMixED does not employ differentially private training algorithms, it does not incur substantial training overhead like DP-SGD. Furthermore, we reduce computational and storage costs by employing parameter-efficient fine-tuning methods. In summary, our key contributions are the following: (1) We introduce and formalize $(\alpha,\beta)$-RD Mollifiers, a generalization of $\epsilon$-Mollifers, to avoid additive DP noise. (2) We propose a private prediction protocol utilizing RD Mollifiers and prove that it satisfies the DP prediction definition with group-level privacy. (3) We experimentally demonstrate that PMixED outperforms DP-SGD on two large-scale datasets. (4) We open-source our software implementation of PMixED to further spark research in this area [^2] https://github.com/james-flemings/pmixed.
@@ 2 Related Works
Most of the previous differential privacy work in LLMs has focused on private training. McMahan et al. (2017) first explored private training of language models using a small recurrent neural network to achieve user-level differential privacy in the federated learning setting. Recent breakthroughs in differentially private LLMs involve self-supervised pre-training on public data, followed by privately fine-tuning on a private corpus Li et al. (2021); Yu et al. (2021).
An orthogonal approach, but conceptually similar to ours, is PATE Papernot et al. (2016, 2018). PATE also uses an ensemble of models trained on pairwise disjoint subsets of a private dataset to generate DP labels. However, PATE and PMixED rely on substantially different private aggregation schemes. PATE uses the Gaussian/Laplacian mechanism to perturb its output vote count histogram while our method utilizes the inherent stochastic nature of sampling from a probability distribution and a public model in LLMs. Furthermore, PATE does not satisfy the DP prediction definition Dwork and Feldman (2018) since its data-dependent privacy accounting causes its privacy loss to be a function of the private data. Hence, testing that the data-dependent loss does not exceed the privacy budget will leak additional privacy Redberg et al. (2023).
@@ 3.2 Differential Privacy
3.3 Private Training vs. Private Prediction
We say that $A$ is a private training algorithm if it produces weights that are differentially private with respect to a private corpus $D$. For DP-SGD, $A(D)$ returns the per-sample gradients of the parameters perturbed with Gaussian noise for each iteration of training. Private training algorithms provide strong privacy since they limit privacy leakage even when an adversary has complete, white-box access to the model parameters.
We observe that commercial deployments of LLMs are typically cloud-based and the model is only accessible through API, so an adversary does not have access to model parameters. In this black box setting an alternative approach, called private prediction Dwork and Feldman (2018), provides DP at prediction, which exploits this important relaxation. We formalize this below:
@@ 5.1 DP Guarantees for PMixED
At a high level, we will prove that $\mathcal{P}$ is $(\alpha,\epsilon_{G}/T)$-RDP for query $\mathbf{x}_{t}$. Then, using the Composition Theorem C.2, we can say that $\mathcal{P}$ is $(\alpha,\epsilon_{G})$-RDP for query-responses $\{(\mathbf{x}_{t},y_{t})\}_{t=1}^{T}$, thus satisfying the Private Prediction Definition 3.4. Appendix B analyzes the case when $\alpha=\infty$, which is pure DP.
@@
where $\lambda_{i}$ is selected from Equation 5. Now, observe that each $\lambda_{i}$ is dependent only on $D_{i}$, so $p_{-i}(\mathbf{x}_{t})$ does not contain $D_{i}$. Using the fact that $D_{\alpha}^{\leftrightarrow}(\overline{p}_{j}(\mathbf{x}_{t})||p_{0}(\mathbf{x}_{t}))\leq\beta\alpha$ for all $j\in[N]$, then for any two neighboring ensembles $\{p_{i}\}_{i=1}^{N}$, $\{p_{j}\}_{j\neq i}$ with $N>1$
@@
by the Quasi Convexity property of Renyi Divergence C.6. Then using the Triangle-like Inquality C.5 gives us $D_{\alpha}(\overline{p}_{i}(\mathbf{x}_{t})||p_{-i}(\mathbf{x}_{t}))\leq 4\beta\alpha$. Hence plugging in Equation 6 for $\beta$ into Equation 8 implies $D_{\alpha}(p(\mathbf{x}_{t})||p_{-i}(\mathbf{x}_{t}))\leq\epsilon_{G}/T$. When $N=1$, then $p_{-i}(\mathbf{x}_{t})=p_{0}(\mathbf{x}_{t})$ since our protocol will resort to using $p_{0}$.
Now, for the other way
$D_{\alpha}(p_{-i}(\mathbf{x}_{t})||p(\mathbf{x}_{t}))$
$=\textstyle D_{\alpha}\left(p_{-i}(\mathbf{x}_{t})\Big{\|}\frac{N-1}{N}p_{-i}(\mathbf{x}_{t})+\frac{1}{N}\overline{p}(\mathbf{x}_{t})\right)$
$\textstyle\leq\frac{N-1}{N}D_{\alpha}(p_{-i}(\mathbf{x}_{t})||p_{-i}(\mathbf{x}_{t}))$
$+\textstyle\frac{1}{N}D_{\alpha}(p_{-i}(\mathbf{x}_{t})||\overline{p}_{i}(\mathbf{x}_{t}))$ (9)
$\leq\frac{4\beta\alpha}{N}$ (10)
$=\frac{\frac{N-1}{N}\log(1)+\frac{1}{N}\log\left(\exp\left(4\beta\alpha(\alpha-1)\right)\right)}{\alpha-1}$
$\leq\frac{\log\left(\frac{N-1+\exp\left({(\alpha-1)4\beta\alpha}\right)}{N}\right)}{\alpha-1}$ (11)
where Equation 9 uses convexity in the second argument of Renyi divergence (Theorem C.7), Equation 10 uses the same argument as in Equation 8, and Equation 11 is due to concavity of logarithms. Then setting $\beta$ to Equation 6 gives us $D_{\alpha}(p_{-i}(\mathbf{x}_{t})||p_{(}\mathbf{x}_{t}))\leq\epsilon_{G}/T$. Thus proves the claim that the output of $\mathcal{P}$ on query $\mathbf{x}_{t}$ is $(\alpha,\epsilon_{G}/T)$-RDP. ∎
@@ 6.1 Experimental Setup
The public model in our experiments is a pre-trained GPT-2 small model. We compare PMixED to three baselines: the public model, a non-private fine-tuned model, and a private fine-tuned model produced by DP-SGD. Although DP-SGD has a weaker privacy level, we compared PMixED to per-sample DP-SGD as our baseline because it illustrates how well PMixED performs against the most widely-used DP solution. The non-private and private fine-tuned baselines also used LoRA. LoRA based fine-tuning can outperform full private fine-tuning since the LoRA updates a much smaller set of parameters during fine-tuningYu et al. (2021). The LoRA parameters used for each model $p_{i}$ contain $0.11\%$ of the total number of parameters of the pre-trained model, allowing us to fit the entire ensemble into one GPU during prediction. Appendix A contains more details about fine-tuning.
@@ 6.3 Ablation Study
7 Conclusion
PMixED is inspired by the observation that most LLM deployments are cloud-based where an adversary only has black-box access to the model, rather than access to the entire model parameters. Under this setting, PMixED presents a novel private prediction protocol that provides DP during prediction, rather than during training. Thus PMixED is model agnostic, which has significant practical implications since it avoids the complexity of privately training models with millions or even billions of parameters. Our approach relies on the novel idea of $(\alpha,\beta)$-RD Mollifiers to project each of the private model’s output distribution onto a set around a public LLM’s output distribution. PMixED achieves the strong group-level DP and outperforms DP-SGD in terms of utility on large-scale datasets using LLMs. Thus, PMixED substantially progresses private prediction methods in LLMs and offers a practical alternative to DP training methods.
@@ 8 Limitations
PMixED performs better with larger ensemble sizes combined with longer training epochs, as quantified in the ablation studies. As ensemble sizes grow each model in the ensemble needs sufficient data to train, which in turn increases the cumulative training data size. We also note that PMixED requires more training epochs than regular SGD. However, these limitations are generally a bottleneck for all DP-based approaches, such as DP-SGD, and hence are not unique limitations of PMixED.
In this work, we employed LoRA to significantly reduce model size demands. However, storing an ensemble of models in memory can place a burden on the computing resources during training and inference procedures. The inference latency can be negatively impacted as an ensemble of models must provide predictions first followed by calculating the lambdas, due to the use of the bisection method. Potential systems optimizations can be to parallelize the inference and lambda calculations. We leave it as a future work to reduce the inference latency. But on the positive side, PMixED operates well with batch-based training and hence can take advantage of parallel GPU resources, while DP-SGD needs per-sample gradients which can curtail GPU efficiency.
Perhaps the biggest limitation of PMixED is the query budget, which limits the number of predictions that can be made. However, this limitation is a natural consequence of differentially private prediction from a nonprivately trained model. As a future work, one can explore a more relaxed problem setup, a PATE-like setting Papernot et al. (2016, 2018), where we produce DP predictions while minimizing the privacy loss.
Finally, we note how work in DP-SGD, as well as this work, uses a pre-trained model to boost performance. Unfortunately, using publicly available data is not necessarily risk-free in terms of privacy, as prior works were able to extract personally identifiable information from a GPT-2 model pre-trained on data scraped from the public internet Carlini et al. (2021). Thus model deployments must still be cautious of using pre-trained models in terms of understanding their information leakage potential.